OpenStudy (anonymous):

The length of a rectangle is 2 more than three times the width. The area of the rectangle is 161 square inches. What are the dimensions of the rectangle? If x = the width of the rectangle, which of the following equations is used in the process of solving this problem?

myininaya (myininaya):

width=x length=l l=2+3x area=A A=161=lx lx=161 (2+3x)x=161 2x+3x^2=161 3x^2+2x-161=0 solve this for x and you will have width then just plug that x into l=2+3x to find length

myininaya (myininaya):

only take positive values of 3x^2+2x-161=0